# On the Real Zeroes of Half-integral Weight Hecke Cusp Forms, II

### v1

@@ Abstract.
We show that for $\gg K^{2}$ of the half-integral weight Hecke cusp forms in the Kohnen plus subspace with weight bounded by a large parameter $K$, the number of ”real” zeroes grows at the expected rate. A key technical step in the proof is to obtain sharp bounds for the mollified first and second moments of quadratic twists of modular $L$-functions.
@@ 1. Introduction and the main result
We are interested in the real [^1] This notion comes from the fact that half-integral weight form is real-value on the lines $\delta_{1}$ and $\delta_{2}$ when it is normalised to have real Fourier coefficients. zeroes of half-integral Hecke cusp forms, that is zeroes on the two geodesic segments
@@
Let us now summarise the main results of [13]. Before that we need to introduce some notation. Throughout the article, let $k$ be a positive integer. We write $S_{k+\frac{1}{2}}(4)$ for the space of half-integral weight cusp forms of weight $k+\frac{1}{2}$ and level $4$. Also we denote by $S_{k+\frac{1}{2}}^{+}(4)\subset S_{k+\frac{1}{2}}(4)$ the Kohnen plus subspace and let $B^{+}_{k+\frac{1}{2}}$ denote a fixed Hecke eigenbasis for $S_{k+\frac{1}{2}}^{+}(4)$. Note that for half-integral weight cusp forms we cannot normalise the coefficient $c_{g}(1)$ to be equal to one without losing the algebraicity of the Fourier coefficients. This means that, unlike in the integral weight case, there is no canonical choice for $B_{k+\frac{1}{2}}^{+}$, but this causes no problems for us. We write $\mathcal{Z}(g):=\{z\in\mathcal{F}:\,g(z)=0\}$ for the set of zeroes of $g\in S_{k+\frac{1}{2}}^{+}(4)$ and let $\mathcal{F}_{Y}:=\{z\in\mathcal{F}\,:\,\textnormal{Im}(z)\geq Y\}$ be a Siegel set, where $\mathcal{F}:=\Gamma_{0}(4)\backslash\mathbb{H}$. Finally, set [^2] Here, and throughout the paper, the notation $\ell\sim L$ means that $L\leq\ell\leq 2L$.
@@
    For $\gg K^{2}/(\log K)^{7/2}$ of the forms $g\in\mathcal{S}_{K}$ we have
@@
In the same article it was alluded that it might be possible to obtain a result that holds for a positive proportion of forms in $\mathcal{S}_{K}$ with $\gg K/Y$ real zeros in $\mathcal{F}_{Y}$ by introducing a suitable mollifier to the argument (we shall review the approach of [13] in the next section). The goal of the present work is to verify this prediction. In the following section we introduce a mollifier that is suitable for our aim. The mollifier $M_{g}(d)$ is chosen so that typically $|c_{g}(|d|)M_{g}(d)|\approx 1$. Our main result is the following, which unifies and improves the results in Theorem 1.1.
@@
In the range $\sqrt{k\log k}\leq Y<\frac{1}{100}k$ it is expected that
@@
This should be compared to a result of Lester, Matomäki, and Radziwiłł [20] for integral weight forms. They showed that for any fixed $\varepsilon>0$ there exists a subset [^3] Throughout the paper we write $H_{k}$ for the set of holomorphic cusp forms of integral weight $k$ and full level. We also write $B_{k}$ for the Hecke eigenbasis of $H_{2k}$. $S_{k}\subset H_{k}$, containing more than $(1-\varepsilon)\#H_{k}$ elements, such that for every $f\in S_{k}$ we have
@@
The proof of their result relies on the breakthrough of Matomäki and Radziwiłł [24] on multiplicative functions in short intervals. Concerning the first result mentioned, note that for individual forms $f$ we cannot do as well, even on the assumption of the Generalised Lindelöf or Generalised Riemann Hypothesis (GRH). The reason is that in order to produce real zeroes of $f$ we look at sign changes of the coefficients $\lambda_{f}(n)$. In order to obtain a positive proportion of the zeroes on the line we need a positive proportion of sign changes between the coefficients of $\lambda_{f}(n)$ in appropriate ranges of $n$. However, we cannot e.g. rule out the scenario where for all $p\leq(\log k)^{2-\varepsilon}$ we have $\lambda_{f}(p)=0$.
@@ 2. The Strategy
(2.1) $\alpha_{g}:=\frac{\Gamma\left(k-\frac{1}{2}\right)}{(4\pi)^{k-\frac{1}{2}}\|g\|_{2}^{2}}>0.$
Throughout this section we set $X=K/Y$. Very briefly, it follows from a steepest descend argument that in order to obtain real zeroes of $g$ one needs to produce fundamental discriminants $d_{\pm}$ in short intervals $x\leq(-1)^{k}d\leq x+H$ with $H$ as small as possible in terms of $X\sim x$ so that $\sqrt{\alpha_{g}}c_{g}(|d_{-}|)<-k^{-\delta}<k^{-\delta}<\sqrt{\alpha_{g}}c_{g}(|d_{+}|)$ for some $\delta>1/2$. For the purpose of exposition we pretend here that we are only looking for numbers $d_{\pm}$ for which $c_{g}(|d_{-}|)<0<c_{g}(|d_{+}|)$. To detect such an event one considers the inequality
@@
We now compare the strategies to compute the mollified moments to the methods used to establish Lemma 2.1 in [13]. Concerning the second moment, the strategy is broadly similar to that in [13]: one executes the $g$-sum using the half-integral weight variant of the Petersson formula after which the $d$- and $k$-sums are evaluated by the Poisson summation. For the mollified second moment we first instead use Waldspurger’s formula to express $|c_{g}(|d|)|^{2}$ as a central $L$-value on which an approximate functional equation is then applied. After that the $d$- and $k$-sums are evaluated by the Poisson summation. The final step is to relate the resulting main term to an Euler product that is estimated by Mertens’ theorem.
For the the unmollified fourth moment was treated by starting with Waldspurger’s formula and an approximate functional equation, after which a large sieve inequality of Deshouillers and Iwaniec was used to estimate the double $(g,k)$-sum. The final step was to average over the fundamental discriminants trivially. With a mollifier this strategy does not work and we need to proceed differently. However, the first steps are identical: by Waldspurger’s formula we need to estimate the sum
@@
where $L(X;d,k)$ is a random $L$-function build from random variables $X$, which model the behaviour of Hecke eigenvalues $\lambda_{f}(n)$. Similarly, $M_{2}(X;d)$ is random mollifier modelling the behaviour of $M_{f}(d)^{4}$. Now, using the Mellin inversion we may express (2.2) as a double line integral
@@ 3. Notations
We use standard asymptotic notation. If $f$ and $g$ are complex-valued functions defined on some set, say $\mathcal{D}$, then we write $f\ll g$ to signify that $|f(x)|\leqslant C|g(x)|$ for all $x\in\mathcal{D}$ for some implicit constant $C\in\mathbb{R}_{+}$. The notation $O(g)$ denotes a quantity that is $\ll g$, and $f\asymp g$ means that both $f\ll g$ and $g\ll f$. We write $f=o(g)$ if $g$ never vanishes in $\mathcal{D}$ and $f(x)/g(x)\longrightarrow 0$ as $x\longrightarrow\infty$. Moreover, we write $f\sim g$ if $f(x)/g(x)\longrightarrow 1$ as $x\longrightarrow\infty$. The letter $\varepsilon$ denotes a positive real number, whose value can be fixed to be arbitrarily small, and whose value can be different in different instances in a proof. All implicit constants are allowed to depend on $\varepsilon$, on the implicit constants appearing in the assumptions of theorem statements, and on anything that has been fixed. When necessary, we will use subscripts $\ll_{\alpha,\beta,...},O_{\alpha,\beta,...}$, etc. to indicate when implicit constants are allowed to depend on quantities $\alpha,\beta,$...
We define $\chi_{d}(\cdot):=\left(\frac{d}{\cdot}\right)$, the Jacobi symbol, for all non-zero odd integers $d$. For such $d$ we set
@@ 4.1. Heuristics
Let $d$ denote a fundamental discriminant, and $\chi_{d}(\cdot)=\left(\frac{d}{\cdot}\right)$ denote the primitive quadratic character of conductor $|d|$. Let $f\otimes\chi_{d}$ denote the twist of $f\in H_{2k}$ by the character $\chi_{d}$, and let $L(s,f\otimes\chi_{d})$ denote the twisted $L$-function
@@
The idea that allows us to take $x$ to be a small power of $X$ is to use an iterative scheme of Radziwiłł and Soundararajan [25], which is partly motivated by the Brun–Hooley sieve. Heuristically we expect that $P(f,d)$ has a Gaussian limiting distribution over $d\sim x$ with mean zero and variance $\sim\log\log x$. We consider the disjoint sets of primes in the intervals $I_{0},I_{1},...,I_{J}$ with $I_{j}=[K^{\theta_{j-1}},K^{\theta_{j}}[$ for j=1,,Jj=1,...,J and set $I_{0}:=]c,K^{\theta_{0}}]$, where $c\geq 2$ is fixed. Due to the expected Gaussian distribution, choosing $\theta_{j}$’s appropriately, we anticipate that each of the terms
@@
$\exp\left(-\frac{1}{2}\sum_{p\in I_{j}}\frac{a_{f}(p)\chi_{d}(p)}{\sqrt{p}}\right)$ $\approx\sum_{0\leq j\leq 10\ell_{j}}\frac{1}{j!}\left(-\frac{1}{2}\sum_{p\in I_{j}}\frac{a_{f}(p)\chi_{d}(p)}{\sqrt{p}}\right)^{j}$
@@
where $\eta_{1}>0$ is sufficiently small absolute constant. We choose $J$ so that $\eta_{2}\leq\theta_{J}\leq e\eta_{2}$ for sufficiently small constant $\eta_{2}>0$.
@@
This mollifier is designed so that it behaves like $L(1/2,f\otimes\chi_{d})^{-1/2}$ for typical $f\in H_{2k}$ and fundamental discriminant $d$.
@@ 4.2. Evaluating the fourth moment
(4.4) $M_{g}(d)^{2\ell}=(\log K)^{\ell/2}\sum_{n\leq K^{2\ell\theta_{J}}}\frac{h_{\ell}(n)a_{f}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}.$
@@
After opening $M_{g}(d)^{4}$ using (4.4), in order to prove Proposition 4.2 we would require an asymptotic evaluation of the multiple average
@@ 5.1. Half-integral weight forms
The Kohnen plus subspace $S_{k+\frac{1}{2}}^{+}(4)\subset S_{k+\frac{1}{2}}(4)$ consists of all weight $k+\frac{1}{2}$ Hecke cusp forms whose $n^{\text{th}}$ Fourier coefficient vanishes whenever $(-1)^{k}n\equiv 2,3\,(\text{mod }4)$. This space has a basis consisting of simultaneous eigenfunctions of the $T(p^{2})$ for odd $p$. It is well-known that, as $k\longrightarrow\infty$, asymptotically one third of half-integral weight cusp forms lie in the Kohnen plus space.
@@
Let $f$ be a Hecke cusp form of weight $2k$ for the full modular group $\mathrm{SL}_{2}(\mathbb{Z})$. Then
@@
where, for any $\sigma>2$,
(5.5) $V_{k}(x):=\frac{1}{2\pi i}\int\limits_{(\sigma)}g(s)x^{-s}e^{s^{2}}\frac{\mathrm{d}s}{s}\quad\text{with}\quad g(s):=(2\pi)^{-s}\frac{\Gamma(s+k)}{\Gamma(k)}.$
Furthermore, for $|\textnormal{Im}(s)|\leq\sqrt{k}$ we have for any $A\geq 1$ that
(5.6) $V_{k}(\xi)=1+O(\xi^{1/2-o(1)}),\qquad V_{k}(\xi)=1+O_{A}(\xi^{-A}).$
@@ 5.2.1. Summation formulas
Main properties of these integral transforms have been worked out by Khan. He has shown [16, (2.17)] that
@@
(5.9) $\frac{\partial^{j}}{\partial\xi^{j}}W_{K}^{(2)}\left(\frac{\xi}{|d|},\frac{z}{|d|},\frac{K^{2}c}{8\pi\sqrt{\xi z}}\right)\ll_{j,A,B}\left(\frac{\xi}{|d|}\right)^{-j-A}\left(\frac{\sqrt{\xi z}}{K^{2}c}\right)^{B}K^{\varepsilon}$
for any $j\geq 0$, $A,B>0$.
@@
$\sum_{k\equiv 0\,(2)}h\left(\frac{k-1}{K}\right)=\frac{K}{2}\widehat{h}(0)+O_{B}(K^{-B})$

### v2

@@ Abstract.
We show that for $\gg K^{2}$ of the half-integral weight Hecke cusp forms in the Kohnen plus subspaces with weight bounded by a large parameter $K$, the number of ”real” zeroes grows at the expected rate. A key technical step in the proof is to obtain sharp bounds for the mollified first and second moments of quadratic twists of modular $L$-functions.
@@ 1. Introduction and the main result
In this paper we are interested in the real [^1] This notion comes from the fact that half-integral weight form is real-value on the lines $\delta_{1}$ and $\delta_{2}$ when it is normalised to have real Fourier coefficients. zeroes of half-integral weight Hecke cusp forms, that is zeroes on the two geodesic segments
@@
Let us now summarise the main results of [13]. Before that we need to introduce some notation. Throughout the article, let $k$ be a positive integer. We write $S_{k+\frac{1}{2}}(4)$ for the space of cusp forms of weight $k+\frac{1}{2}$ and level $4$. Also we denote by $S_{k+\frac{1}{2}}^{+}(4)\subset S_{k+\frac{1}{2}}(4)$ the Kohnen plus subspace and let $B^{+}_{k+\frac{1}{2}}$ denote a fixed Hecke eigenbasis for $S_{k+\frac{1}{2}}^{+}(4)$. Note that for half-integral weight cusp forms we cannot normalise the coefficient $c_{g}(1)$ to be equal to one without losing the algebraicity of the Fourier coefficients. This means that, unlike in the integral weight case, there is no canonical choice for $B_{k+\frac{1}{2}}^{+}$, but this causes no problems for us. We write $\mathcal{Z}(g):=\{z\in\mathcal{F}:\,g(z)=0\}$ for the set of zeroes of $g\in S_{k+\frac{1}{2}}^{+}(4)$ and let $\mathcal{F}_{Y}:=\{z\in\mathcal{F}\,:\,\textnormal{Im}(z)\geq Y\}$ be a Siegel set, where $\mathcal{F}:=\Gamma_{0}(4)\backslash\mathbb{H}$. Finally, set [^2] Here, and throughout the paper, the notation $\ell\sim L$ means that $L\leq\ell\leq 2L$.
@@
    For $\gg_{\varepsilon}K^{2}/(\log K)^{3/2+\varepsilon}$ of the forms $g\in\mathcal{S}_{K}$ we have
@@
In the same article it was alluded that it might be possible to obtain a result that holds for a positive proportion of forms in $\mathcal{S}_{K}$ with $\gg K/Y$ real zeroes in $\mathcal{F}_{Y}$ by introducing a suitable mollifier to the argument (we shall review the approach of [13] in the next section). The goal of the present work is to verify this prediction. In the following section we introduce a mollifier that is suitable for our aim. The mollifier $M_{g}(d)$ is chosen so that typically $|c_{g}(|d|)M_{g}(d)|\approx 1$. Our main result is the following, which unifies and improves the results in Theorem 1.1.
@@
(1) It might be possible to extract an explicit proportion of forms for which Theorem 1.2 holds from our argument, but this seems difficult.
(2) In the range $\sqrt{k\log k}\leq Y<\frac{1}{100}k$ it is expected that
@@
Theorem 1.2 should be compared to a result of Lester, Matomäki, and Radziwiłł [20] for integral weight forms. They showed that for any fixed $\varepsilon>0$ there exists a subset [^3] Throughout the paper we write $H_{k}$ for the set of holomorphic cusp forms of integral weight $k$ and full level. We also write $B_{k}$ for the Hecke eigenbasis of $H_{2k}$. $S_{k}\subset B_{k}$, containing more than $(1-\varepsilon)\#B_{k}$ elements, such that for every $f\in S_{k}$ we have
@@
The proof of their result relies on the breakthrough of Matomäki and Radziwiłł [24] on multiplicative functions in short intervals. Concerning the first result mentioned, note that for individual forms $f$ we cannot do as well, even on the assumption of the Generalised Lindelöf or Generalised Riemann Hypothesis (GRH). The reason is that in order to produce real zeroes of $f$ we look at sign changes of the Hecke eigenvalues $\lambda_{f}(n)$. In order to obtain a positive proportion of the zeroes on the line we need a positive proportion of sign changes between the coefficients of $\lambda_{f}(n)$ in appropriate ranges of $n$. However, we cannot e.g. rule out the scenario where for all $p\leq(\log k)^{2-\varepsilon}$ we have $\lambda_{f}(p)=0$.
@@ 2. The Strategy
(2.1) $\alpha_{g}:=\frac{\Gamma\left(k-\frac{1}{2}\right)}{2(4\pi)^{k-\frac{1}{2}}\|g\|_{2}^{2}}>0.$
Throughout this section we set $X=K/Y$. Very briefly, it follows from a steepest descend argument that in order to obtain real zeroes of $g$ one needs to produce fundamental discriminants $d_{\pm}$ in short intervals $x\leq(-1)^{k}d\leq x+H$ with $H$ as small as possible in terms of $X\sim x$ so that $\sqrt{\alpha_{g}}c_{g}(|d_{-}|)<-k^{-\delta}<k^{-\delta}<\sqrt{\alpha_{g}}c_{g}(|d_{+}|)$ for some sufficiently small $\delta>1/2$. For the purpose of exposition we pretend here that we are only looking for numbers $d_{\pm}$ for which $c_{g}(|d_{-}|)<0<c_{g}(|d_{+}|)$. To detect such an event one considers the inequality
@@
We now compare the strategies to compute the mollified moments to the methods used to establish Lemma 2.1 in [13]. Concerning the second moment, in [13] one executes the $g$-sum using the half-integral weight variant of the Petersson trace formula after which the $d$- and $k$-sums are evaluated by the Poisson summation. For the mollified second moment we first instead use Waldspurger’s formula to express $|c_{g}(|d|)|^{2}$ as a central $L$-value on which an approximate functional equation is then applied. After that the $d$- and $k$-sums are evaluated by the Poisson summation. The final step is to relate the resulting main term to an Euler product that is estimated by Mertens’ theorem.
The unmollified fourth moment was treated by starting with Waldspurger’s formula and an approximate functional equation, after which a large sieve inequality of Deshouillers and Iwaniec was used to estimate the double $(g,k)$-sum. The final step was to average over the fundamental discriminants trivially. With a mollifier this strategy does not work and we need to proceed differently. However, the first steps are identical: by Waldspurger’s formula we need to estimate the sum
@@
where $L(X;d,k)$ is a random $L$-function build from random variables $X$, which model the behaviour of Hecke eigenvalues $\lambda_{f}(n)$. Similarly, $M_{2}(X;d)$ is random mollifier modelling the behaviour of $M_{f}(d)^{4}$. Now, using Mellin inversion we may express (2.2) as a double line integral
@@ 3. Notations
We use standard asymptotic notation. If $f$ and $g$ are complex-valued functions defined on some set, say $\mathcal{G}$, then we write $f\ll g$ to signify that $|f(x)|\leqslant C|g(x)|$ for all $x\in\mathcal{G}$ for some implicit constant $C\in\mathbb{R}_{+}$. The notation $O(g)$ denotes a quantity that is $\ll g$, and $f\asymp g$ means that both $f\ll g$ and $g\ll f$. We write $f=o(g)$ if $g$ never vanishes in $\mathcal{G}$ and $f(x)/g(x)\longrightarrow 0$ as $x\longrightarrow\infty$. Moreover, we write $f\sim g$ if $f(x)/g(x)\longrightarrow 1$ as $x\longrightarrow\infty$. The letter $\varepsilon$ denotes a positive real number, whose value can be fixed to be arbitrarily small, and whose value can be different in different instances in a proof. All implicit constants are allowed to depend on $\varepsilon$, on the implicit constants appearing in the assumptions of theorem statements, and on anything that has been fixed. When necessary, we will use subscripts $\ll_{\alpha,\beta,...},O_{\alpha,\beta,...}$, etc. to indicate when implicit constants are allowed to depend on quantities $\alpha,\beta,$...
We define $\chi_{d}(\cdot):=\left(\frac{d}{\cdot}\right)$, the Kronecker symbol, for all non-zero odd integers $d$. For such $d$ we set
@@ 4.1. Heuristics
Let $d$ denote a fundamental discriminant, and $\chi_{d}(\cdot)=\left(\frac{d}{\cdot}\right)$ denote the primitive quadratic character of conductor $|d|$. Let $f\otimes\chi_{d}$ denote the twist of $f\in B_{k}$ by the character $\chi_{d}$, and let $L(s,f\otimes\chi_{d})$ denote the twisted $L$-function
@@
The idea that allows us to take $x$ to be a small power of $X$ is to use an iterative scheme of Radziwiłł and Soundararajan [25], which is partly motivated by the Brun–Hooley sieve. Heuristically we expect that $P(f,d)$ has a Gaussian limiting distribution over $d\sim x$ with mean zero and variance $\sim\log\log x$. We consider the disjoint sets of primes in the intervals $I_{0},I_{1},...,I_{J}$ with $I_{j}=[K^{\theta_{j-1}},K^{\theta_{j}}[$ for j=1,,Jj=1,...,J and set $I_{0}:=]c,K^{\theta_{0}}]$, where $c>2$ is fixed and chosen to be sufficiently large. Due to the expected Gaussian distribution, choosing $\theta_{j}$’s appropriately, we anticipate that each of the terms
@@
$\exp\left(-\frac{1}{2}\sum_{p\in I_{j}}\frac{a_{f}(p)\chi_{d}(p)}{\sqrt{p}}\right)$ $\approx\sum_{0\leq\ell\leq 10\ell_{j}}\frac{1}{\ell!}\left(-\frac{1}{2}\sum_{p\in I_{j}}\frac{a_{f}(p)\chi_{d}(p)}{\sqrt{p}}\right)^{\ell}$
@@
where $\eta_{1}>0$ is sufficiently large absolute constant. We choose $J$ so that $\eta_{2}\leq\theta_{J}\leq e\eta_{2}$ for sufficiently small constant $\eta_{2}>0$.
@@
This mollifier is designed so that it behaves like $L(1/2,f\otimes\chi_{d})^{-1/2}$ for typical $f\in B_{k}$ and fundamental discriminant $d$. Note that as $c>2$ it follows that $M_{g}(d;j)$ is supported only on odd $n$’s. Throughout the paper we set
$\delta_{0}:=\sum_{j=0}^{J}\ell_{j}\theta_{j}$
and note that $M_{g}(d)\ll K^{\delta_{0}}(\log K)^{1/4}$. By choosing $\eta_{2}$ to be small we can guarantee that $\delta_{0}$ is small enough. This will be important in Section $12$.
@@ 4.2. Evaluating the fourth moment
(4.4) $M_{g}(d)^{2\ell}=(\log K)^{\ell/2}\sum_{n\leq K^{2\ell\delta_{0}}}\frac{h_{\ell}(n)a_{f}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}.$
@@
After opening $M_{g}(d)^{4}$ using (4.4) and using Waldspurger’s formula to express $|c_{g}(|d|)|^{2}$ as a central value of an $L$-function, in order to prove Proposition 4.2 we would require an asymptotic evaluation of the multiple average
@@ 5.1. Half-integral weight forms
The Kohnen plus subspace $S_{k+\frac{1}{2}}^{+}(4)\subset S_{k+\frac{1}{2}}(4)$ consists of all weight $k+\frac{1}{2}$ Hecke cusp forms whose $n^{\text{th}}$ Fourier coefficient vanishes whenever $(-1)^{k}n\equiv 2,3\,(\text{mod }4)$. This space has a basis consisting of simultaneous eigenfunctions of $T(p^{2})$ for odd $p$. It is well-known that, as $k\longrightarrow\infty$, asymptotically one third of half-integral weight cusp forms lie in the Kohnen plus space.
@@
Let $f$ be a Hecke cusp form of weight $2k$ for the full modular group $\mathrm{SL}_{2}(\mathbb{Z})$. Then for any fundamental discriminant $d$ for which $(-1)^{k}d>0$ we have that
@@
where, for any $\sigma>1/2$,
(5.5) $V_{k}(x):=\frac{1}{2\pi i}\int\limits_{(\sigma)}\eta(s)x^{-s}e^{s^{2}}\frac{\mathrm{d}s}{s}\quad\text{with}\quad\eta(s):=(2\pi)^{-s}\frac{\Gamma(s+k)}{\Gamma(k)}.$
Furthermore, we have that
(5.6) $V_{k}(\xi)=1+O\left(\left(\frac{\xi}{k}\right)^{1/2-o(1)}\right)$
as $\xi\longrightarrow 0$.
@@ 5.2.1. Summation formulas
Main properties of these integral transforms have been worked out by Khan [^5] In [16] the definition of $V_{k}(\xi)$ differs slightly from ours, but identical arguments to those in [16] lead to same bounds for expressions involving $V_{k}(\xi)$ in our situation.. He has shown [16, (2.17)] that
@@
(5.9) $\frac{\partial^{j}}{\partial\xi^{j}}W_{K}^{(2)}\left(\frac{\xi}{|d|},\frac{z}{|d|},\frac{K^{2}c}{8\pi\sqrt{\xi z}}\right)\ll_{j,A_{1},A_{2},B}\left(\frac{|d|K}{\xi}\right)^{A_{1}}\xi^{-j}\left(\frac{|d|K}{z}\right)^{A_{2}}\left(\frac{\sqrt{\xi z}}{K^{2}c}\right)^{B}K^{\varepsilon}$
for any $j\geq 0$, $A_{1},A_{2},B>0$.
@@
$\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)=\frac{K}{2}\widehat{h}(0)+O_{B}(K^{-B})$